SUB-MILLIMETER WAVE SPECTROSCOPY OF HNC PRODUCED IN AN EXTENDED NEGATIVE GLOW DISCHARGE

\underline{T. AMANO}, M. NEZU, AND Z. ZELINGER^{a}, Institute for Astrophysics and Planetary Sciences, Ibaraki University, Mito, Japan 310-8512.

We demonstrated that HNC produced in an extended negative glow discharge from a gas mixture of CH_{4} and N_{2} exhibited the same magnetic field dependence as ionic species, and we concluded that HNC was produced predominantly through a dissociative recombination reaction of HCNH^{+} with electrons^{b}. Okabayashi and Tanimoto^{c} observed the transitions up to J = 4 − 3 for all the vibrational fundamentals as well as for the ground state. They found that the vibrational temperature was 230 K for the ν_{2} state, while those for the ν_{1} and ν_{3} modes were 1300-1600 K in a glow discharge in a mixture of CH_{3}CN and He. In our extended negative glow discharge, the vibrational temperature was found to be about 1500 K with no distinction among the vibrational modes. We extended the measurements to higher-J(up to J = 7 ← 6) and higher vibrational excited states using frequency-magnetic field double modulation technique. An extensive analysis was performed, in particular, for higher bending vibrational states. We will discuss implication of our observation to the branching ratio of the dissociative recombination reaction of HCNH^{+}.

aJapan Space Forum Research Associate

bT. Amano, paper RD05, 53rd International Symposium on Molecular Spectroscopy, Columbus, Ohio, USA, June, 1998

cT. Okabayashi and M. Tanimoto, J. Chem. Phys. 99, 3268-3271(1993)